43. Justin Industries produces three versions of tires: Supreme, Advanced, and Basic. A condensed segmented income statementJustin Industries produces three versions of tires: Supreme, Advanced, and Basic. A condensed segmented income statement for a recent period follows:Assume all of the fixed expenses for the hard rubber line are avoidable. What will be total net income if the line is dropped?A) $250,000B)$206,000C)$210,000D) $280,00044. Which of the following costs are variable?A) 1 and 2B)1 and 4C)only 1D) only 2Page 10SupremeAdvancedBasicTotalSales$1,000,000$400,000$130,000$1,530,000Variable expenses 650,000 280,000 116,000 1,046,000Contribution margin350,000120,00014,000484,000Fixed expenses 150,000 70,000 44,000 264,000Net income (loss)$200,000$ 50,000$(30,000)$220,000
Answer #1

1) Solution: $250,000

Explanation: When the line is dropped their will no loss of $30,000 thus it will be added back in the net income of $220,000. Thus total net income when the line is dropped equals $220,000 + $30,000 = $250,000

2) Solution: 1 and 4

1st option: 10,000/100,000=0.1; 30,000/300,000 = 0.1

4th option: 10000/50000 = 0.2; 30000/150000=.2

1st option and 4th option gives the same per unit cost thus are variable costs